OpenStudy (anonymous):

complete the square 2x^2+8x+1=0

OpenStudy (anonymous):

\[2x^2+8x=-1\] \[x^2+4x=-\frac{1}{2}\] \[(x+2)^2=-\frac{1}{2}+4=\frac{7}{2}\] \[x+2=\pm\sqrt{\frac{7}{2}}\] done in one more step